OpenStudy (anonymous):

Part 1: Write the equation of the line that passes through the points (3, 4) and (–3, –8) in point-slope form. Part 2: Using complete sentences, explain whether or not it matters which point is used in the final answer. Also explain why you chose the point you did.

OpenStudy (anonymous):

y=mx+b m=4-(-8) ------ 3-(-3) m=12/6=2 y=2x+b 4=2(3)+b 4=6+b -2=b y=2x-2

OpenStudy (anonymous):

part 2. No it doesnt matter which point u use because u would get the same answer for both since they lie on the same line. i choose (3,4) because they were positive numbers